# What the flock knows that the birds do not: exploring the emergence of joint agency in multi-agent active inference

### v1

@@ 1 Introduction
The fundamental unit of analysis in biology and cognition is often the agent—an entity, such as a person, animal, or even a single neuron or cell, that possesses a well-defined boundary and some degree of agency, that is, autonomy in perceiving and interacting with an environment. However, many of the most intriguing phenomena in biological and cognitive systems emerge not from isolated agents but from the collective dynamics of multiple interacting ones. Examples include the coordinated behaviors and self-organization of cells composing a body morphology (levin2019computational; friston2015knowing; pezzulo2016top; levin2017endogenous; manicka2025field)neural population dynamics underlying brain function (freeman1994characterization; fuchs1992phase; friston2012perception; deco2015rethinking; hesse2014self), the collective behaviors animals (holldobler2009superorganism; krause2002living; sumpter2010collective), distributed models of mind (minsky1986society; hofstadter1999godel), and even human societies viewed as collective agents (sawyer2005social). Such collective intelligence phenomena occur across multiple scales of biological and cognitive organization (mcmillen2024collective).
@@
Notions of joint agency appear at many levels of organization and cognitive sophistication. At the level of human social behavior, classical theories of social cognition emphasize both social interaction and collective behavior: while individuals act autonomously, they also participate in shared social structures. Theories of mentalizing highlight that agents form beliefs about, and models of, other agents’ minds (frith2005theory). Complementary accounts emphasize shared and aligned representations, in which multiple agents maintain overlapping internal models of their environment (pezzulo2025predictive). In these frameworks, joint agency corresponds to phenomena such as shared representations (sebanz2006joint; Sebanz2021), collective intentions and goals (Bratman2013; butterfill2023towards; tomasello2005understanding), joint commitments (gilbert2020social), common ground (Clark1996), joint payoffs and team reasoning (Sugden2003), and “we-representations” of action and intention (Gallotti2013). Empirical work in cognitive neuroscience supports these views, identifying mechanisms such as neural coupling (Hasson2012; Hasson2016; Marsh2009; Keller2016), mirroring (Gallese2004; rizzolatti2004mirror) and sensorimotor communication (Pezzulo2019) as enabling the fine-grained spatiotemporal coordination required for joint action. Subjectively, such coordination can give rise to a shared sense of agency—the feeling that outcomes were caused together (pacherie2014does).
@@
Computational models have begun to formalize these social dynamics in terms of interacting agents that communicate, coordinate, and collaborate toward shared objectives. These frameworks typically assume internal (generative) models that encode distinctions between self and others—“my,” “your,” and “our” actions and intentions (Wolpert2003; kilner2007mirror). More recent approaches, however, explore interactive dynamics in which multiple agents maintain and update a shared world model to minimize prediction errors toward a common goal (Friston2015; Maisto2023; Friston2024). Related concepts include the notion of agent-neutral models, or internal models that predict the collective consequences of joint actions regardless of who executed them (pezzulo2017avoiding; pezzulo2025predictive), and shared beliefs such as “public beliefs” (Foerster2019) or the “imagined we” (Tang2022), where collective cognitive dimensions (e.g., beliefs, plans, agency)—shared across multiple individuals—supersede and drive individual cognitive states.
@@
Here we are concerned with a more primitive notion of joint agency that arises at lower levels of organization: at the level of collective behavior and self-organization among simple, particle-like agents that lack advanced cognitive abilities and rich internal models incorporating notions like “our beliefs” or “our plans”. This primitive form of joint agency emerges from the simple fact that teams of agents can infer their own states and actions based on those of other, surrounding agents to which they are informationally coupled—and they are capable of making collective decisions (couzin2005effective; sridhar2021geometry). There is a long tradition of studying the self-organization of collectives, such as active particles, animal swarms, and robot ensembles, using methods from statistical physics and information theory (dorigo2021swarm; bialek2012statistical; cavagna2010scale; bechinger2016active; gomez2023fish; abdel2022self). Complementary approaches have also been developed to quantify causal emergence and the extent to which higher-level collective dynamics exhibit causal power beyond that of their individual components (hoel2013quantifying; rosas2020reconciling).
@@
An emerging trend is the study of collective phenomena and multi-agent systems within the active inference framework (Parr2022; Friston2015; Maisto2023; heins2024collective; levchuk2019active; beckenbauer2025orchestrator; ruiz2024factorised). Active inference was initially developed to address the cognitive and neural processes associated with isolated biological organisms and their action-perception cycle. The general idea was that a single imperative or objective function — the minimization of variational free energy — suffices to explain both perception and action and their associated neural dynamics in biological organisms.
@@
Recent developments have extended active inference principles to the collective dynamics of multiple agents, simpler (e.g., active particles) or more complex. The main difference between active inference and typical statistical approaches used to study collective phenomena is that each component is a full-fledged agent, with its action-perception cycle and minimizing its free energy based on local signals from the environment and/or other agents. For instance, in a network of agents playing the role of “neurons,” each minimizing its own variational free energy, a single neuron can infer whether or not to fire based on the activity of surrounding neurons. The ensemble of neurons can thereby exhibit synchronous dynamics and become collectively responsive to perceptual stimuli and reward contingencies (palacios2019emergence; gandolfi2022emergence; gandolfi2025network). Similarly, in a collection of cells engaged in morphogenesis and pattern formation, each cell can infer its own position in the final body morphology from the chemical signals emitted by neighboring cells, while simultaneously emitting signals that guide others (friston2015knowing). Functionally, each cell acts as an individual active inference agent minimizing its own free energy, yet its collective dynamics lead to the emergence of a coherent body morphology that is resilient—for example, capable of reconstructing itself after perturbation.
@@
When endowed with joint agency, active inference agents do not lose their individual autonomy: each continues to infer and minimize its own free energy. However, their collective agency at the system level supersedes individual agency to some extent, as the fate of each agent becomes jointly determined by the states and signals of others (heins2024collective). This formulation allows studying agency at two (or more) nested levels: the level of the single agent and the agent collective (and of a collective composed of collectives, and so on). The formulation also allows studying how collective processes can go awry, for example, when a single agent becomes insensitive to signals from other agents and hence myopically pursues individual rather than collective goals. In this perspective, a breakdown of cell-cell communication that causes individual cells to prioritize unicellular objectives rather than large-scale, collective morphogenetic goals has been proposed as a possible mechanism for cancer (levin2021bioelectric; levin2021bioelectrical).
@@
Despite significant progress in modeling the collective behavior of multiple (simple) active inference agents, the relationships between these models and the broader notions of joint agency remain only partially understood. Previous active inference simulations have primarily focused on the self-organization of agents into cohesive multi-agent structures—such as bodies or coordinated ensembles—but have paid less attention to the functional consequences of this self-organization for higher-order phenomena such as joint agency or collective beliefs. These functional notions are typically not explicitly encoded in the internal generative models of the simulated agents, unlike in active inference models of higher-level cognition (e.g., human–human joint action (Maisto2023)), where such constructs are explicitly represented.
@@ 2 Formalizing individual and joint agency in flocking behavior through Markov blankets
We develop a simulation of flocking behavior, in which we consider an ensemble of 100 active inference agents (”birds”), each endowed with identical internal models and each minimizing its local variational free energy. To infer its heading direction, each bird uses observations about the heading directions of its neighbors. This simple mechanism promotes the self-organization of the birds into a collective flocking behavior. It is analogous to models in statistical physics (wu1982potts) and flocking simulations in computer graphics (reynolds1987flocks), but it is based on local inference rather than on predefined rules (see Section A for a formal specification of the active inference agents).
Figure 1: Formalizing individual and joint agency in flocking behavior through Markov blankets (a) Example simulation of flocking behavior among 100 birds over 8 time steps, showing the gradual alignment of headings. (b) Markov blankets, individual and joint agency. Schematic representation of an individual agent, whose internal states are separated from the external environment by a Markov blanket (dashed line). The blanket mediates the exchange of information through sensory $s$ and active $a$ states, defining the agent’s boundary for perception and action. (c) Illustration of a group of interacting agents whose collective dynamics are enclosed within a higher-level Markov blanket (dashed line). Through reciprocal coupling and shared information flow, the ensemble functions as a single, higher-order agent (’flock’), exemplifying the emergence of joint or shared agency from multiple interacting components. (d) The emergent Markov blanket around the flock during time steps 1–6 of the simulation in (a). Colors illustrate the statistical roles of individual birds relative to the flock Markov Blanket: internal (blue), active (red), and sensory (magenta). External states are depicted as cyan birds.
@@
To assess whether this alignment can be formalized as a transition toward joint agency—namely, from individual birds to a flock—we turn to the notion of (nested) Markov blankets (kirchhoff2018markov; parr2020markov; friston2023path). A Markov blanket defines a statistical boundary around an agent, delineating the interface between its internal states and the external environment—and thereby its domain of perception and action (pearl2014probabilistic) (Figure 1B). It is based on the notion of statistical independence: internal and external states are conditionally independent given a set of blanket states, typically partitioned into sensory and active states. Sensory states mediate the flow of information from the environment to the internal states (observations), while active states mediate the influence of internal dynamics on the environment (actions). This statistical separation enables an agent to maintain its identity and autonomy by inferring hidden causes in the environment from its sensory inputs, and by acting to minimize the discrepancy between predicted and observed states.
Crucially, the Markov blanket formalism allows for hierarchical nesting of agents and collectives. When multiple agents become sufficiently coupled through reciprocal sensory and active exchanges, their collective dynamics can give rise to a higher-level Markov blanket encompassing the group as a whole (Figure 1C). In this configuration, some of the individual agents’ states function as the internal, sensory, and active components of a superordinate blanket, thereby defining a new, emergent agent. Thus, while each agent retains its individual agency, the group simultaneously acquires a higher-level, joint agency. Such nesting provides a principled, statistical account of how joint agency can arise naturally from the interactions of individual agents: as informational boundaries reorganize, new levels of autonomy and coordinated inference emerge (friston2013life; heins2024collective).
@@ 3 Sensing and escaping predators in the flock
Figure 2B shows the energy of the system, defined as in the vector Potts model with four states and Moore neighborhood (wu1982potts), which characterizes the average degree of alignment among the birds (and thus approximates the size or coherence of the flock). When the predator appears at an early stage (red dotted line), the energy continues to increase briefly before decreasing, indicating that disorganization occurs with some delay. In contrast, when the predator appears later (blue dotted line), the energy is initially much higher due to the presence of an already well-formed flock; here, the decrease in energy occurs more rapidly, reflecting a faster predator-induced destabilization.
Figure 2C shows the dynamics of the “stress” state across birds, which—as expected—increase sharply when the predator appears and then gradually decline. This pattern is expected, since the stress propagation is designed to be independent of the agents’ current states.
@@ 4 Synergistic information about predator location in the flock
To quantify this, we employ partial information decomposition (PID) (williams2010nonnegative), an information-theoretic framework that measures how two source variables (e.g., a pair of birds) together provide information about a target variable (e.g., the location of the predator). Crucially, PID decomposes total mutual information into three components: the unique information provided by each source individually, the redundant information shared between them, and the synergistic information that emerges only when both sources are considered jointly. Synergistic information has been proposed as a foundation for formalizing emergence, capturing higher-level informational structure beyond the sum of individual contributions (rosas2020reconciling).
@@
Figure 3 summarizes these results. The core concepts of PID, including unique information, redundancy, and synergy, are conceptually illustrated in Figure 3A. Figure 3B summarizes the temporal dynamics of all four PID atoms, averaged across all bird pairs. Notably, the unique information atoms (top panels) and the redundancy atom (bottom left panel) remain comparable to zero, falling within the gray band of the null distribution at all time steps. In sharp contrast, the synergy atom (bottom right panel) is the only variable that shows significant values. Specifically, we observe a significant peak in synergistic information for the second predator appearance (blue curve). In contrast, the synergy related to the first predator (red curve) shows a similar temporal shape (a peak, decline, and rebound), but it falls within the null distribution band and is therefore not statistically significant. The significant synergy (blue curve) peaks shortly after the predator’s appearance and then declines, with a smaller rebound thereafter.
@@
Summing up, this analysis reveals that the flock exhibits synergistic information about the predator’s location, and that this information is stronger and more spatially distributed during the second predator event, when a larger flock is present. This finding illustrates a form of collective knowledge that extends beyond the information available to individual birds, and may help explain the faster collective response observed during the second predator encounter. It also points toward a notion of causal emergence, supported by the presence of synergistic information (rosas2020reconciling).

### v2

@@ 1 Introduction
The fundamental unit of analysis in biology and cognition is often the agent—an entity, such as a person, animal, or even a single neuron or cell, that possesses a well-defined boundary and some degree of agency, that is, autonomy in perceiving and interacting with an environment. However, many of the most intriguing phenomena in biological and cognitive systems emerge not from isolated agents but from the collective dynamics of multiple interacting ones. Examples include the coordinated behaviors and self-organization of cells composing a body morphology (Levin, 2019; Friston et al., 2015; Pezzulo and Levin, 2016; Levin et al., 2017; Manicka and Levin, 2025)neural population dynamics underlying brain function (Freeman, 1994; Fuchs et al., 1992; Friston et al., 2012; Deco et al., 2015; Hesse and Gross, 2014), the collective behaviors animals (Hölldobler and Wilson, 2009; Krause and Ruxton, 2002; Sumpter, 2010), distributed models of mind (Minsky, 1986; Hofstadter, 1999), and even human societies viewed as collective agents (Sawyer, 2005). Such collective intelligence phenomena occur across multiple scales of biological and cognitive organization (McMillen and Levin, 2024).
@@
Notions of joint agency appear at many levels of organization and cognitive sophistication. At the level of human social behavior, classical theories of social cognition emphasize both social interaction and collective behavior: while individuals act autonomously, they also participate in shared social structures. Theories of mentalizing highlight that agents form beliefs about, and models of, other agents’ minds (Frith and Frith, 2005). Complementary accounts emphasize shared and aligned representations, in which multiple agents maintain overlapping internal models of their environment (Pezzulo et al., 2025a). In these frameworks, joint agency corresponds to phenomena such as shared representations (Sebanz et al., 2006; Sebanz and Knoblich, 2021), collective intentions and goals (Bratman, 2013; Butterfill and Sinigaglia, 2023; Tomasello et al., 2005), joint commitments (Gilbert, 2020), common ground (Clark, 1996), joint payoffs and team reasoning (Sugden, 2003), and “we-representations” of action and intention (Gallotti and Frith, 2013). Empirical work in cognitive neuroscience supports these views, identifying mechanisms such as neural coupling (Hasson et al., 2012; Hasson and Frith, 2016; Marsh et al., 2009; Keller et al., 2016), mirroring (Gallese et al., 2004; Rizzolatti and Craighero, 2004) and sensorimotor communication (Pezzulo et al., 2019) as enabling the fine-grained spatiotemporal coordination required for joint action. Subjectively, such coordination can give rise to a shared sense of agency—the feeling that outcomes were caused together (Pacherie, 2014).
@@
Computational models have begun to formalize these social dynamics in terms of interacting agents that communicate, coordinate, and collaborate toward shared objectives. These frameworks typically assume internal (generative) models that encode distinctions between self and others—“my,” “your,” and “our” actions and intentions (Wolpert et al., 2003; Kilner et al., 2007). More recent approaches, however, explore interactive dynamics in which multiple agents maintain and update a shared world model to minimize prediction errors toward a common goal (Friston and Frith, 2015; Maisto et al., 2023; Friston et al., 2024). Related concepts include the notion of agent-neutral models, or internal models that predict the collective consequences of joint actions regardless of who executed them (Pezzulo et al., 2017, 2025a), and shared beliefs such as “public beliefs” (Foerster et al., 2019) or the “imagined we” (Tang et al., 2022), where collective cognitive dimensions (e.g., beliefs, plans, agency)—shared across multiple individuals—supersede and drive individual cognitive states.
@@
Here we are concerned with a more primitive notion of joint agency that arises at lower levels of organization: at the level of collective behavior and self-organization among simple, particle-like agents that lack advanced cognitive abilities and rich internal models incorporating notions like “our beliefs” or “our plans”. This primitive form of joint agency emerges from the simple fact that teams of agents can infer their own states and actions based on those of other, surrounding agents to which they are informationally coupled—and they are capable of making collective decisions (Couzin et al., 2005; Sridhar et al., 2021). There is a long tradition of studying the self-organization of collectives, such as active particles, animal swarms, and robot ensembles, using methods from statistical physics and information theory (Dorigo et al., 2021; Bialek et al., 2012; Cavagna et al., 2010; Bechinger et al., 2016; Gómez-Nava et al., 2023; Abdel-Rahman et al., 2022; Durve et al., 2020). Complementary approaches have also been developed to quantify causal emergence and the extent to which higher-level collective dynamics exhibit causal power beyond that of their individual components (Hoel et al., 2013; Rosas et al., 2020).
@@
An emerging trend is the study of collective phenomena and multi-agent systems within the active inference framework (Parr et al., 2022; Friston and Frith, 2015; Maisto et al., 2023; Heins et al., 2024; Levchuk et al., 2019; Beckenbauer et al., 2025; Ruiz-Serra et al., 2024). Active inference was initially developed to address the cognitive and neural processes associated with isolated biological organisms and their action-perception cycle. The general idea was that a single imperative or objective function — the minimization of variational free energy — suffices to explain both perception and action and their associated neural dynamics in biological organisms.
@@
Recent developments have extended active inference principles to the collective dynamics of multiple agents, simpler (e.g., active particles) or more complex. The main difference between active inference and typical statistical approaches used to study collective phenomena is that each component is a full-fledged agent, with its action-perception cycle and minimizing its free energy based on local signals from the environment and/or other agents. For instance, in a network of agents playing the role of “neurons,” each minimizing its own variational free energy, a single neuron can infer whether or not to fire based on the activity of surrounding neurons. The ensemble of neurons can thereby exhibit synchronous dynamics and become collectively responsive to perceptual stimuli and reward contingencies (Palacios et al., 2019; Gandolfi et al., 2022, 2025). Similarly, in a collection of cells engaged in morphogenesis and pattern formation, each cell can infer its own position in the final body morphology from the chemical signals emitted by neighboring cells, while simultaneously emitting signals that guide others (Friston et al., 2015). Functionally, each cell acts as an individual active inference agent minimizing its own free energy, yet its collective dynamics lead to the emergence of a coherent body morphology that is resilient—for example, capable of reconstructing itself after perturbation.
@@
When endowed with joint agency, active inference agents do not lose their individual autonomy: each continues to infer and minimize its own free energy. However, their collective agency at the system level supersedes individual agency to some extent, as the fate of each agent becomes jointly determined by the states and signals of others (Heins et al., 2024). This formulation allows studying agency at two (or more) nested levels: the level of the single agent and the agent collective (and of a collective composed of collectives, and so on). The formulation also allows studying how collective processes can go awry, for example, when a single agent becomes insensitive to signals from other agents and hence myopically pursues individual rather than collective goals. In this perspective, a breakdown of cell-cell communication that causes individual cells to prioritize unicellular objectives rather than large-scale, collective morphogenetic goals has been proposed as a possible mechanism for cancer (Levin, 2021a, b).
@@
Despite significant progress in modeling the collective behavior of multiple (simple) active inference agents, the relationships between these models and the broader notions of joint agency remain only partially understood. Previous active inference simulations have primarily focused on the self-organization of agents into cohesive multi-agent structures—such as bodies or coordinated ensembles—but have paid less attention to the functional consequences of this self-organization for higher-order phenomena such as joint agency or collective beliefs. These functional notions are typically not explicitly encoded in the internal generative models of the simulated agents, unlike in active inference models of higher-level cognition (e.g., human–human joint action (Maisto et al., 2023)), where such constructs are explicitly represented.
@@ 2 Formalizing individual and joint agency in flocking behavior through Markov blankets
We develop a simulation of flocking behavior, in which we consider an ensemble of 100 active inference agents (”birds”), each endowed with identical internal models and each minimizing its local variational free energy. To infer its heading direction, each bird uses observations about the heading directions of its neighbors. This simple mechanism promotes the self-organization of the birds into a collective flocking behavior. It is analogous to models in statistical physics (Wu, 1982) and flocking simulations in computer graphics (Reynolds, 1987), but it is based on local inference rather than on predefined rules (see Section A for a formal specification of the active inference agents).
Figure 1: Formalizing individual and joint agency in flocking behavior through Markov blankets (a) Example simulation of flocking behavior among 100 birds over 8 time steps, showing the gradual alignment of headings. (b) Markov blankets, individual and joint agency. Schematic representation of an individual agent, whose internal states are separated from the external environment by a Markov blanket (dashed line). The blanket mediates the exchange of information through sensory $s$ and active $a$ states, defining the agent’s boundary for perception and action. (c) Illustration of a group of interacting agents whose collective dynamics are enclosed within a higher-level Markov blanket (dashed line). Through reciprocal coupling and shared information flow, the ensemble functions as a single, higher-order agent (’flock’), exemplifying the emergence of joint or shared agency from multiple interacting components. Colors illustrate the statistical roles of individual birds relative to the flock Markov Blanket: internal (blue), active (red), and sensory (magenta). External states are depicted as cyan birds. Note that all the birds functioning as internal states are oriented in the same direction, whereas this coordination is not required for birds functioning as sensory and active states. (d) The emergent Markov blanket around the flock during time steps 1–6 of the simulation in (a). The figure illustrates an average over the first 6 timesteps.
@@
To assess whether this alignment can be formalized as a transition toward joint agency—namely, from individual birds to a flock—we turn to the notion of (nested) Markov blankets (Kirchhoff et al., 2018; Parr et al., 2020; Friston et al., 2023). A Markov blanket defines a statistical boundary around an agent, delineating the interface between its internal states and the external environment—and thereby its domain of perception and action (Pearl, 2014) (Figure 1B). It is based on the notion of statistical independence: internal and external states are conditionally independent given a set of blanket states, typically partitioned into sensory and active states. Sensory states mediate the flow of information from the environment to the internal states (observations), while active states mediate the influence of internal dynamics on the environment (actions). This statistical separation enables an agent to maintain its identity and autonomy by inferring hidden causes in the environment from its sensory inputs, and by acting to minimize the discrepancy between predicted and observed states.
Crucially, the Markov blanket formalism allows for hierarchical nesting of agents and collectives. When multiple agents become sufficiently coupled through reciprocal sensory and active exchanges, their collective dynamics can give rise to a higher-level Markov blanket encompassing the group as a whole (Figure 1C). In this configuration, some of the individual agents’ states function as the internal, sensory, and active components of a superordinate blanket, thereby defining a new, emergent agent. Thus, while each agent retains its individual agency, the group simultaneously acquires a higher-level, joint agency. Such nesting provides a principled, statistical account of how joint agency can arise naturally from the interactions of individual agents: as informational boundaries reorganize, new levels of autonomy and coordinated inference emerge (Friston, 2013; Heins et al., 2024).
@@ 3 Sensing and escaping predators in the flock
Figure 2B shows the energy of the system, defined as in the vector Potts model with four states and Moore neighborhood (see Equation 1 in A.2) (Wu, 1982), which characterizes the average degree of alignment among the birds (and thus approximates the size or coherence of the flock). When the predator appears at an early stage (red dotted line), the energy continues to increase briefly before decreasing, indicating that disorganization occurs with some delay. In contrast, when the predator appears later (blue dotted line), the energy is initially much higher due to the presence of an already well-formed flock; here, the decrease in energy occurs more rapidly, reflecting a faster predator-induced destabilization.
Figure 2C shows the dynamics of the grand mean (i.e., the mean over the simulations of means over birds) of the “stress” states (introduced in A.2) across birds, which—as expected—increase sharply when the predator appears and then gradually decline. This pattern is expected, since the stress propagation is designed to be independent of the agents’ current states.
@@ 4 Synergistic information about predator location in the flock
To quantify this, we employ partial information decomposition (PID) (Williams and Beer, 2010), an information-theoretic framework that measures how two source variables (e.g., a pair of birds) together provide information about a target variable (e.g., the location of the predator). Crucially, PID decomposes total mutual information into three components: the unique information provided by each source individually, the redundant information shared between them, and the synergistic information that emerges only when both sources are considered jointly. Synergistic information has been proposed as a foundation for formalizing emergence, capturing higher-level informational structure beyond the sum of individual contributions (Rosas et al., 2020).
@@
Figure 3 summarizes these results. The core concepts of PID, including unique information, redundancy, and synergy, are formally introduced in A.5 and conceptually illustrated in Figure 3A. Figure 3B summarizes the temporal dynamics of all four PID atoms, averaged across all bird pairs. Notably, the unique information atoms (top panels) and the redundancy atom (bottom left panel) remain comparable to zero, falling within the gray band of the null distribution at all time steps. In sharp contrast, the synergy atom (bottom right panel) is the only variable that shows significant values. Specifically, we observe a significant peak in synergistic information for the second predator appearance (blue curve). In contrast, the synergy related to the first predator (red curve) shows a similar temporal shape (a peak, decline, and rebound), but it falls within the null distribution band and is therefore not statistically significant. The significant synergy (blue curve) peaks shortly after the predator’s appearance and then declines, with a smaller rebound thereafter.
@@
Summing up, this analysis reveals that the flock exhibits synergistic information about the predator’s location, and that this information is stronger and more spatially distributed during the second predator event, when a larger flock is present. This finding illustrates a form of collective knowledge that extends beyond the information available to individual birds, and may help explain the faster collective response observed during the second predator encounter. It also points toward a notion of causal emergence, supported by the presence of synergistic information (Rosas et al., 2020).